16 May 2010

Sunny Sunday

Gas lab under construction
Alli and Theo mounting shelves in the new gas laboratory.

Things progress at a tremendous rate. However, we cannot begin core processing before we have all equipment, and the remaining equipment is on the next plane, which is now scheduled for Tuesday. The bad weather last week did not only hit us, but also other ice sheet sites, so the 109th are behind schedule. But today the weather was fine, and people enjoyed being on the ice sheet.

The GRIT traverse left today at Noon, we all said good-bye to the four participants. They now head for Summit Camp. It was indeed a pleasure to have them visit, and we wish them a safe journey.

What we have done today:
1. Drilled. One run yielding a 3.45 m core.
2. Unpacking science equipment.
3. The new warm laboratory is now fully installed with tables and shelves.
4. Working on insulating the ventilation funnels of CFA and gas laboratories.
5. Fixing small bugs in the drillers program.
6. Logging ice cores and discussing procedures.
7. Cleaning up in the cargo line.
8. Japanese pit-study in progress.
9. Setting up water-vapour sampling site.

Ad.1: Drillers made one run where they drilled a nice 3.45 m long core.
The run was normal, and the drillers, Trevor and Steff, had drillers classes. An extra support for the tower has been made.

Ad 6: We have agreed to process the core this year in 1.10 m  sections.
Thus the loggers will make a cut every 1.10 m. In each trough in the core buffer, there is room for 2 x 1.10 m. We can thus use the core buffer to full capacity.

Weather: blue sky, -28°C to - 18°C, 5-9 knots from SE. Visibility: Unrestricted.

FL, J.P. Steffensen


← Previous entry   Next entry →